George Kyparisis owns a company that manufactures sailboats. Actual demand for George's sailboats during each of the past four seasons was as follows: 

Season1234
Winter1,4401,2401,000960
Spring1,5001,4401,6401,540
Summer1,0002,1002,0402,000
Fall640750670520

George has forecasted that annual demand for his sailboats in year 5 will equal 6,500 sailboats. 

Based on the given data and using the multiplicative seasonal model, the demand level for George's sailboats in the spring of year 5 will be _______  sailboats (enter a whole number).
Answer #1

1. SPRING = 1814

EXPLANATION:

SEASON

1

2

3

4

WINTER

1440

1200

1040

960

SPRING

1560

1400

1620

1540

SUMMER

1000

2120

2020

1900

FALL

600

750

650

560

AVERAGE

1150

1367.5

1332.5

1240

AVERAGE 1 = 1440 + 1560 + 1000 + 600 / 4 = 1150

AVERAGE 2 = 1200 + 1400 + 2120 + 750 / 4 = 1367.5

AVERAGE 3 = 1040 + 1620 + 2020 + 650 / 4 = 1332.5

AVERAGE 4 = 960 + 1540 + 1900 + 560 / 4 = 1240



SEASON

1

2

3

4

AVERAGE SI

WINTER

1440 / 1150 = 1.25217391304348

1200 / 1367.5 = 0.877513711151737

1040 / 1332.5 = 0.780487804878049

960 / 1240 = 0.774193548387097

1.252173913 + 0.8775137112 + 0.7804878049 + 0.7741935484 / 4 = 0.9211

SPRING

1560 / 1150 = 1.35652173913043

1400 / 1367.5 = 1.02376599634369

1620 / 1332.5 = 1.21575984990619

1540 / 1240 = 1.24193548387097

1.356521739 + 1.023765996 + 1.21575985 + 1.241935484 / 4 = 1.2095

SUMMER

1000 / 1150 = 0.869565217391304

2120 / 1367.5 = 1.55027422303473

2020 / 1332.5 = 1.51594746716698

1900 / 1240 = 1.53225806451613

0.8695652174 + 1.550274223 + 1.515947467 + 1.532258065 / 4 = 1.367

FALL

600 / 1150 = 0.521739130434783

750 / 1367.5 = 0.548446069469835

650 / 1332.5 = 0.487804878048781

560 / 1240 = 0.451612903225806

0.5217391304 + 0.5484460695 + 0.487804878 + 0.4516129032 / 4 = 0.5024

ANNUAL SALES NEXT YEAR = 6000

AVERAGE SALES PER SEASON = 1500

ADJUSTED DEMAND

SEASON

1

CALCULATION

WINTER

1382

1500 * 0.9211 = 1382

SPRING

1814

1500 * 1.2095 = 1814

SUMMER

2051

1500 * 1.367 = 2051

FALL

754

1500 * 0.5024 = 754